OpenStudy (anonymous):

What is when f(x) = - 2x3 + x2 + 16x - 15 and g(x) = x + 3

OpenStudy (chosenmatt):

what do you think you need to do solve

OpenStudy (anonymous):

im not sure how to solve it tbh

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

@skate-4-life270

OpenStudy (chosenmatt):

@CausticSyndicalist

OpenStudy (anonymous):

what

OpenStudy (chosenmatt):

well i don't get what your question is asking...

OpenStudy (anonymous):

do you need it graphed?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-2x2 - 5x + 1 - 2x2 + 7x - 5 -2x2 + 7x - 5 - - 2x2 - 5x + 1 -

OpenStudy (loser66):

Your question: What is......???? what?? missed this part

OpenStudy (anonymous):

those are the answer choices

OpenStudy (anonymous):

oh what do the two have in common

OpenStudy (anonymous):

What is when f(x) = - 2x3 + x2 + 16x - 15 and g(x) = x + 3?

OpenStudy (anonymous):

what?

OpenStudy (loser66):

Again, it is no meaning!!

OpenStudy (loser66):

what is f(x) * g(x) when.... what is f(g(x) ) when.... what is f/g(x) when.... what is f(x) + g(x) when... which one is your problem?

OpenStudy (anonymous):

yeah man sorry but idk this one

OpenStudy (anonymous):

the question is What is when f(x) = - 2x3 + x2 + 16x - 15 and g(x) = x + 3? but i dont understand how to answer it